OpenStudy (anonymous):

SOMEONE PLEASE HELP ME ON THIS :( An equation is shown below: 6(2x - 11) + 15 = 21 Write the steps you will use to solve the equation and explain each step.

OpenStudy (kitkat16):

first you need to solve for x yes?

sammixboo (sammixboo):

Simplify as much as yo can. The only thing right now you can simplify is 6(2x - 11) . Do you know how to distribute?

OpenStudy (anonymous):

I read the lesson, but I didn't really understand it.

OpenStudy (anonymous):

Can you please explain it to me?

OpenStudy (kitkat16):

multiply 6 by each term 6*2 and 6*11 12x-66+15=21 now add 15 to -66 which = -51 12x-51=21 now add 51 to both sides 12x=51+21 51+21=72 12x=72 now solve for x x= you can either multiply or divide 12*x=72 x=???

OpenStudy (anonymous):

i get it now!!!

OpenStudy (anonymous):

would this be a good way to set up my answer?

OpenStudy (anonymous):

Step 1: 6(2x - 11) + 15 = 21 Step 2: 12x - 66 + 15 = 21 Step 3: 12x - 51 = 21 Step 4: 12x = 72 The answer is: 6

OpenStudy (anonymous):

Do you think that's okay? :)

OpenStudy (kitkat16):

I would add this under step 3 step 4 12x=51+21 step 5 divide - x=72/12 x=6

OpenStudy (kitkat16):

good job!!

OpenStudy (kitkat16):

sorry that isn't -x up there just a dash so ignore

OpenStudy (anonymous):

oh okay, well thank you so so much for all your help beautiful! i appreciate it :)